# Efficient Best-of-Both-Worlds Algorithms for Contextual Combinatorial Semi-Bandits

### v1

@@ 3.1 Numerical Experiments
We now evaluate the per-iteration runtime of Algorithm 3. All experiments are conducted on a machine with a 2.3 GHz 8-core Intel Core i9 processor and all optimization problems are modeled in Python. In all experiments, we fix the $m$-set size to $m=5$, vary the number of base arms $K\in\{10,\dots,100\}$, and run each algorithm for $N=25$ iterations on a loss vector $y\in[-1,1]^{K}$ whose entries are drawn uniformly at random. Mean per-iteration runtimes and their 95% confidence intervals are reported, and two instances of problem (1), each employing a different regularizer, are evaluated. The first instance uses Tsallis entropy with parameter $\alpha=1/2$, as in (zimmert2019beating, ; zimmert2021tsallis, ), which is known to achieve best-of-both-worlds results. This corresponds to the regularizer $f(x)=-\sqrt{x}$ (labeled as “Tsallis”). The second instance employs the widely used negative Shannon entropy, induced by the regularizer $f(x)=x\ln{x}$ (labeled as “Negative Shannon entropy”).
@@
We compute the mean-action projection of FTRL via Algorithm 3, and compare it against two baselines: the heuristic Newton method used in (zimmert2019beating, ),  [^1] The source code for this method is available at https://github.com/diku-dk/CombSemiBandits. and a direct implementation of the optimization step using MOSEK. [^2] MOSEK is accessed via the Python Fusion API mosek , as using a higher-level modeling interface such as CVXPY diamond2016cvxpy —with MOSEK as the backend—would prevent a fair runtime comparison. For convenience, we also provide a CVXPY-based implementation leveraging MOSEK. In all cases, we solve to an error tolerance of $\varepsilon=10^{-9}$, matching the suboptimality and feasibility tolerances used in MOSEK.
Figure 1 visualizes the per-iteration runtimes of Algorithm 3, the Newton method, and MOSEK, as a function of the number $K$ of base arms. We observe that Algorithm 3 runs nearly 10 times faster than the Newton baseline for $K=100$, and consistently outperforms MOSEK by a factor of approximately 20 across all values of $K$. This highlights the computational efficiency of our bisection-based Algorithm 3 for regularizers whose gradients admit closed-form inverses.
@@ 4 Concluding Remarks and Limitations
We introduce the first algorithm for contextual combinatorial semi-bandits that simultaneously guarantees $\widetilde{\mathcal{O}}(\sqrt{T})$ regret in the adversarial regime and $\widetilde{\mathcal{O}}(\ln T)$ in the corrupted stochastic regime. At its core is a Follow-the-Regularized-Leader (FTRL) scheme with a Shannon-entropy regularizer with time-varying learning rate. Vanilla FTRL, however, requires solving a $K$-dimensional convex projection in each round, which limits scalability, whereas Follow-the-Perturbed-Leader (FTPL) is favored for its speed. We recover FTPL-style speed-ups by exploiting the Karush-Kuhn-Tucker (KKT) conditions to reduce the projection to a one-dimensional root-finding problem. This hybrid design preserves FTRL’s strong theoretical guarantees while matching FTPL’s per-round efficiency, enabling practical implementations that achieve substantial per-iteration speed-ups. Despite these advances, our approach has two main limitations. First, the Shannon-entropy regularizer introduces an additional $\widetilde{\mathcal{O}}(\ln T)$ term in the adversarial regret. Second, the dependence on the combinatorial action size $m$ scales as $\sqrt{m}$, which is suboptimal relative to known lower bounds (zierahn2023nonstochastic, ). Closing the $\sqrt{m}$ gap, extending the fast-projection technique to other regularizers are promising directions for further enhancing both the theory and practice of best-of-both-worlds bandit algorithms.
Acknowledgement
This work was supported as a part of the NCCR Automation, a National Center of Competence in Research, funded by the Swiss National Science Foundation (grant number 51NF40_225155).
@@ A.1 Proofs and Auxiliary Results for Section 2
For ease of notations in the proofs, we set $\gamma_{t}=\alpha_{t}\eta_{t}$ throughout this section. We start by addressing the terms relevant to the regret induced in the auxiliary game for a fixed context $\widetilde{\mathcal{R}}_{T}(x).$ Denote as well the time-varying Bregman divergence from $p\in\operatorname*{conv}(\mathcal{A})$ to $q\in\operatorname*{conv}(\mathcal{A})$ through
$D_{t}(q,p)=\psi_{t}(q)-\psi_{t}(p)-\langle\nabla\psi_{t}(q),q-p\rangle.$
@@
$\mathbb{E}_{A_{t}}\left[\sum_{t=1}^{T}\langle x,\tilde{\theta}_{t,k}\rangle((A_{t})_{k}-(u^{\star}(x))_{k})\right]$ (2)
@@
$\mathbb{E}_{A_{t}}$ $\left[\sum_{t=1}^{T}\langle x,\tilde{\theta}_{t,k}\rangle\big{(}(A_{t})_{k}-(u^{\star}(x))_{k}\big{)}\right]$
@@
Applying a standard FTRL regret decomposition result (see e.g. [21, Exercise 28.12]) to the first term on the right-hand side of the above expression and applying the definition of $U(x)$ yields the desired result. ∎
@@
where the second inequality exploits Hölder’s inequality, which applies because $\|X_{t}\|_{2}\leq 1$ by assumption, the third inequality holds thanks to [38, Lemma 1], the fourth inequality holds due to the parameter choice $\gamma_{t}=\alpha_{t}\eta_{t}$. The last inequality again follows by the parameter choices $\alpha_{t}=4K\ln(t)/(\gamma_{t}\lambda_{\min}(\Sigma))$ and $\eta_{t}\leq 1/2$. ∎
@@
We analyze the stability and penalty terms appearing in Lemma A.1. First, we bound the per-round stability term in (2). By [21, Theorem 26.13] and using that $\tfrac{\partial^{2}}{(\partial x)^{2}}\psi_{t}(x)=\tfrac{1}{\eta_{t}x}$ for all $x\in\operatorname*{conv}(\mathcal{A})$, we have
@@
$\sum_{k=1}^{K}\langle x,\tilde{\theta}_{t,k}\rangle^{2}(z_{t})_{k}\leq 2\sum_{k=1}^{K}\langle x,\tilde{\theta}_{t,k}\rangle^{2}(\bar{A}_{t}(x))_{k}.$ (4)
Substituting (4) into (3) establishes the upper bound for the stability term as claimed. As for the penalty term, we have
@@
[38, Lemma 3] and our choice of $E$ that satisfies $|E|=K$ imply that
$\max_{A\in\mathcal{A}}\mathbb{E}\Bigg{[}\sum_{k=1}^{K}\langle x,\hat{\theta}_{t,k}-\tilde{\theta}_{t,k}\rangle(A)_{k}\mid\mathcal{F}_{t-1}\Bigg{]}\leq\sqrt{m}\exp\left(-\frac{\gamma_{t}\lambda_{\min}(\Sigma)M_{t}}{2K}\right).$
@@
The following is another preparation lemma for establishing Lemma 2.4, which constitutes an adaptation of [20, Lemma 18] to the combinatorial semi-bandit setting. It provides a refined bound on the penalty term induced by the Shannon entropy regularizer with respect to the ghost sample.
@@
$=2c_{1}\sqrt{m\ln(K/m)}\mathbb{E}\left[\sum_{t=1}^{T}\frac{H(\bar{A}_{t+1}(X_{0}))(\sqrt{\sum_{s=1}^{t+1}H(\bar{A}_{s}(X_{s}))}-\sqrt{\sum_{s=1}^{t}H(\bar{A}_{s}(X_{s}))})}{H(\bar{A}_{t+1}(X_{t+1}))}\right]$
@@
Lemma A.6 (Variance control [38, Lemma 5]).
@@
where the equality holds due to the law of iterated expectations, the first inequality holds thanks to Lemma A.6 and (5), and the second inequality follows from Jensen’s inequality.
@@
$=\mathcal{O}\left(\frac{K(\ln T)^{2}}{c_{1}\lambda_{\min}(\Sigma)\sqrt{\ln(K/m)}}\sqrt{\sum_{t=1}^{T}\mathbb{E}[H(A_{t}(X_{t}))}]\right),$
where the second inequality holds because $\gamma_{t}=\alpha_{t}\eta_{t}$ and the choice of $\alpha_{t}$ as well as the fact that $\ln t\leq\ln T$. The equality then holds because of (5).
@@
where the first inequality is due to the fact that $\beta_{t}=\beta_{t+1}$ while $t\in\left[t_{0}-2\right],\beta_{t_{0}-1}\geq\beta_{t_{0}-1}^{\prime}$ by construction, and $\beta_{t}^{\prime}=\beta_{t}$ for $t\geq t_{0}$. The second inequality holds because $\beta^{\prime}_{t}$ is increasing in $t$, while the second equality holds thanks to Lemma A.5. The claim then follows by substituting (6), (7), and (8) into terms in the statement of Lemma A.3 (to bound the regret for the auxiliary game) as well as Lemma 2.2 (to bound the regret for the original game). ∎
@@
Observe that for any $A\in\operatorname*{conv}(\mathcal{A})$ and any $S\subset\mathcal{A}$ with $|S|=m$, it follows that
@@
$\leq\sum_{k\notin S}(A)_{k}\ln\frac{K-m}{\sum_{k\notin S}(A)_{k}}+\sum_{k\in S}(A)_{k}\left(\frac{1}{(A)_{k}}-1\right)$
$=\sum_{k\notin S}(A)_{k}\left(\ln\frac{K-m}{\sum_{k\notin S}(A)_{k}}+1\right),$
where the inequality holds because of Jensen’s inequality and because $\ln(1/x)\leq 1/x-1$, and the second equality holds because $\sum_{k\in S}(A)_{k}+\sum_{k\notin S}(A)_{k}=m$ thanks to the membership of $A$ in $\operatorname*{conv}(\mathcal{A})$. Applying the above inequality to the entropy with respect to action given the ghost sample $X_{0}$ and $S=\{k\in[K]:(u^{\star}(X_{0}))_{k}=1\}$ gives
$\begin{split}\sum_{t=1}^{T}H(\bar{A}_{t}(X_{0}))&\leq\sum_{t=1}^{T}\sum_{k:(u^{\star}(X_{0}))_{k}=0}(\bar{A}_{t})_{k}\left(\ln\frac{K-m}{\sum_{k:(u^{\star}(X_{0}))_{k}=0}(\bar{A}_{t})_{k}}+1\right)\\ &\leq\sum_{t=1}^{T}\sum_{k:(u^{\star}(X_{0}))_{k}=0}(\bar{A}_{t})_{k}\ln\frac{e(K-m)T}{\sum_{t=1}^{T}\sum_{k:(u^{\star}(X_{0}))_{k}=0}(\bar{A}_{t})_{k}}\\ &\leq\ln((K-m)T)\sum_{t=1}^{T}\sum_{k:(u^{\star}(X_{0}))_{k}=0}(\bar{A}_{t})_{k},\end{split}$ (9)
@@
Substituting the above expression into (9) and noticing that $\mathbb{E}[\pi_{t}(u^{\star}(X_{0}))|\mathcal{F}_{t-1}]=\mathbb{E}[\pi_{t}(u^{\star}(X_{t}))|\mathcal{F}_{t-1}]$ thus establishes the claim. ∎
@@
The regret bound stated in (i) follows immediately from Lemma 2.2 combined with Lemma 2.4 as well as the observation that $\sum_{t=1}^{T}H(\bar{A}_{t}(X_{0}))\leq m\ln(K/m)T.$ We now show claim (ii). Note that in the case of $\sum_{t=1}^{T}\sum_{k:(u^{\star}(X_{0}))_{k}=0}(\bar{A}_{t})_{k}<e$, we have $\sum_{t=1}^{T}H(\bar{A}_{t}(X_{0}))\leq e\ln(e(K-m)T)+1/e$, in which case we already have the desired bound. We thus continue to consider the case when $\sum_{t=1}^{T}\sum_{k:(u^{\star}(X_{0}))_{k}=0}(\bar{A}_{t})_{k}\geq e$. Observe that due to the definition of suboptimality gap $\Delta_{A}(x)$ and the corruption budget $C$, it follows that
@@
where the first inequality holds thanks to Lemma 2.2 combined with Lemma 2.4 as well as the lower bound (10), the second inequality follows by Lemma 2.5. The third inequality holds by the observation $a\sqrt{x}-bx\leq a^{2}/(4b)$ for any nonnegative scalars $a,b,x$, and the last equality follows by choosing $\lambda=1.$

### v2

@@ 3.1 Numerical Experiments
We now evaluate the per-iteration runtime of Algorithm 3. All experiments are conducted on a machine with a 2.3 GHz 8-core Intel Core i9 processor and all optimization problems are modeled in Python. In all experiments, we fix the $m$-set size to $m=5$, vary the number of base arms $K\in\{10,\dots,100\}$, and run each algorithm for $N=25$ iterations on a loss vector $y\in[0,1]^{K}$ whose entries are drawn uniformly at random. Mean per-iteration runtimes and their 95% confidence intervals are reported, and two instances of problem equation 1, each employing a different regularizer, are evaluated. The first instance uses Tsallis entropy with parameter $\alpha=1/2$, as in (Zimmert et al., 2019; Zimmert and Seldin, 2021), which is known to achieve best-of-both-worlds results. This corresponds to the regularizer $f(x)=-\sqrt{x}$ (labeled as “Tsallis”). The second instance employs the widely used negative Shannon entropy, induced by the regularizer $f(x)=x\ln{x}$ (labeled as “Negative Shannon entropy”).
@@
We compute the mean-action projection of FTRL via Algorithm 3, and compare it against two baselines: the heuristic Newton method used in (Zimmert et al., 2019),  [^2] The source code for this method is available at https://github.com/diku-dk/CombSemiBandits. and a direct implementation of the optimization step using MOSEK. [^3] MOSEK is accessed via the Python Fusion API ApS (2025). In all cases, we solve to an error tolerance of $\varepsilon=10^{-7}$, matching the suboptimality and feasibility tolerances used in MOSEK. Figure 1 visualizes the per-iteration runtimes of Algorithm 3, the Newton method, and MOSEK, as a function of the number $K$ of base arms. We observe that Algorithm 3 runs nearly 10 times faster than the Newton baseline for $K=100$, and consistently outperforms MOSEK by a factor of approximately 5 across all values of $K$. This highlights the computational efficiency of our bisection-based Algorithm 3.
@@ 4 Concluding Remarks and Limitations
We introduce the first algorithm for contextual combinatorial semi-bandits that simultaneously guarantees $\widetilde{\mathcal{O}}(\sqrt{T})$ regret in the adversarial regime and $\widetilde{\mathcal{O}}(\ln T)$ in the corrupted stochastic regime. At its core is a Follow-the-Regularized-Leader (FTRL) scheme with a Shannon-entropy regularizer with time-varying learning rate. Vanilla FTRL, however, requires solving a $K$-dimensional convex projection in each round, which limits scalability, whereas Follow-the-Perturbed-Leader (FTPL) is favored for its speed. We recover FTPL-style speed-ups by exploiting the KKT conditions to reduce the projection to a one-dimensional root-finding problem. This hybrid design preserves FTRL’s strong theoretical guarantees while matching FTPL’s per-round efficiency. Despite these advances, our approach has two main limitations. First, the Shannon-entropy regularizer introduces an additional $\widetilde{\mathcal{O}}(\ln T)$ term in the adversarial regret. Second, the dependence on the combinatorial action size $m$ scales as $\sqrt{m}$, which is suboptimal relative to known lower bounds (Zierahn et al., 2023). Addressing the aforementioned limitations would be a promising direction for further enhancing both the theory and practice of best-of-both-worlds bandit algorithms.
Acknowledgements
This work was supported as part of the NCCR Automation, a National Center of Competence in Research, funded by the Swiss National Science Foundation (grant number 51NF40_225155).
@@ A.1 Proofs and Auxiliary Results for Section 2
For ease of notation in the proofs, we set $\gamma_{t}=\alpha_{t}\eta_{t}$ throughout this section. We start by addressing the terms relevant to the regret induced in the auxiliary game for a fixed context $\widetilde{\mathcal{R}}_{T}(x).$ Denote as well the time-varying Bregman divergence from $p\in\operatorname*{conv}(\mathcal{A})$ to $q\in\operatorname*{conv}(\mathcal{A})$ through
$D_{t}(q,p)=\psi_{t}(q)-\psi_{t}(p)-\langle\nabla\psi_{t}(p),q-p\rangle.$
@@
$\mathbb{E}_{A_{t}}\left[\sum_{t=1}^{T}\sum_{k=1}^{K}\langle x,\tilde{\theta}_{t,k}\rangle((A_{t})_{k}-(u^{\star}(x))_{k})\right]$ (2)
@@
$\mathbb{E}_{A_{t}}$ $\left[\sum_{t=1}^{T}\sum_{k=1}^{K}\langle x,\tilde{\theta}_{t,k}\rangle\big((A_{t})_{k}-(u^{\star}(x))_{k}\big)\right]$
@@
Applying a standard FTRL regret decomposition result (see e.g. (Lattimore and Szepesvári, 2020, Exercise 28.12)) to the first term on the right-hand side of the above expression and applying the definition of $U(x)$ yields the desired result. ∎
@@
where the second inequality exploits Hölder’s inequality, which applies because $\|X_{t}\|_{2}\leq 1$ by assumption, the third inequality holds thanks to (Zierahn et al., 2023, Lemma 1), the fourth inequality holds due to the parameter choice $\gamma_{t}=\alpha_{t}\eta_{t}$. The last inequality again follows by the parameter choices $\alpha_{t}=4K\ln(t)/\lambda_{\min}(\Sigma)$ and $\eta_{t}\leq 1/2$. ∎
@@
We analyze the stability and penalty terms appearing in Lemma A.1. First, we bound the per-round stability term in equation 2. By (Lattimore and Szepesvári, 2020, Theorem 26.13) and using that $\tfrac{\partial^{2}}{(\partial x)^{2}}\psi_{t}(x)=\tfrac{1}{\eta_{t}x}$ for all $x\in\operatorname*{conv}(\mathcal{A})$, we have
@@
$\sum_{k=1}^{K}\langle x,\tilde{\theta}_{t,k}\rangle^{2}(z_{t})_{k}\leq e\sum_{k=1}^{K}\langle x,\tilde{\theta}_{t,k}\rangle^{2}(\bar{A}_{t}(x))_{k}.$ (4)
Substituting equation 4 into equation 3 establishes the upper bound for the stability term as claimed. As for the penalty term, we have
@@
(Zierahn et al., 2023, Lemma 3) and our choice of $E$ that satisfies $|E|=K$ imply that
$\max_{A\in\mathcal{A}}\mathbb{E}\Bigg[\sum_{k=1}^{K}\langle x,\hat{\theta}_{t,k}-\tilde{\theta}_{t,k}\rangle(A)_{k}\mid\mathcal{F}_{t-1}\Bigg]\leq\sqrt{m}\exp\left(-\frac{\gamma_{t}\lambda_{\min}(\Sigma)M_{t}}{2K}\right)\quad\forall x\in\mathcal{X}.$
@@
The following is another preparation lemma for establishing Lemma 2.4, which constitutes an adaptation of (Kuroki et al., 2024, Lemma 18) to the combinatorial semi-bandit setting. It provides a refined bound on the penalty term induced by the Shannon entropy regularizer with respect to the ghost sample.
@@
$=2c_{1}\sqrt{m\ln(K/m)}\mathbb{E}\left[\sum_{t=1}^{T}\frac{H(\bar{A}_{t+1}(X_{0}))(\sqrt{\sum_{s=1}^{t+1}H(\bar{A}_{s}(X_{s}))}-\sqrt{\sum_{s=1}^{t}H(\bar{A}_{s}(X_{s}))})}{H(\bar{A}_{t+1}(X_{0}))}\right]$
@@
Lemma A.6 (Variance control (Zierahn et al., 2023, Lemma 5)).
@@
where the equality holds due to the law of iterated expectations, the first inequality holds thanks to Lemma A.6 and equation 5, and the second inequality follows from Jensen’s inequality.
@@
$=\mathcal{O}\left(\frac{K(\ln T)^{2}}{c_{1}\lambda_{\min}(\Sigma)\sqrt{\ln(K/m)}}\sqrt{\sum_{t=1}^{T}\mathbb{E}[H(\bar{A}_{t}(X_{t}))]}\right),$
where the second inequality holds because $\gamma_{t}=\alpha_{t}\eta_{t}$ and the choice of $\alpha_{t}$ as well as the fact that $\ln t\leq\ln T$. The equality then holds because of equation 5.
@@
where the first inequality is due to the fact that $\beta_{t}=\beta_{t+1}$ while $t\in\left[t_{0}-2\right],\beta_{t_{0}-1}\geq\beta_{t_{0}-1}^{\prime}$ by construction, and $\beta_{t}^{\prime}=\beta_{t}$ for $t\geq t_{0}$. The second inequality holds because $\beta^{\prime}_{t}$ is increasing in $t$, while the second equality holds thanks to Lemma A.5. The claim then follows by substituting equation 6, equation 7, and equation 8 into terms in the statement of Lemma A.3 (to bound the regret for the auxiliary game) as well as Lemma 2.2 (to bound the regret for the original game). ∎
@@
Observe that for any $A\in\operatorname*{conv}(\mathcal{A})$ and any $S\subset[K]$ with $|S|\leq m$, it follows that
@@
$\leq\sum_{k\notin S}(A)_{k}\ln\frac{K-|S|}{\sum_{k\notin S}(A)_{k}}+\sum_{k\in S}(A)_{k}\left(\frac{1}{(A)_{k}}-1\right)$
$=\sum_{k\notin S}(A)_{k}\left(\ln\frac{K-|S|}{\sum_{k\notin S}(A)_{k}}+1\right),$
where the inequality holds because of Jensen’s inequality and because $\ln(1/x)\leq 1/x-1$, and the second equality holds because $\sum_{k\in S}(A)_{k}+\sum_{k\notin S}(A)_{k}=|S|$ thanks to the membership of $A$ in $\operatorname*{conv}(\mathcal{A})$. Applying the above inequality to the entropy with respect to action given the ghost sample $X_{0}$ and $S=\{k\in[K]:(u^{\star}(X_{0}))_{k}=1\}$ gives
$\begin{split}\sum_{t=1}^{T}H(\bar{A}_{t}(X_{0}))&\leq\sum_{t=1}^{T}\sum_{k:(u^{\star}(X_{0}))_{k}=0}(\bar{A}_{t})_{k}\left(\ln\frac{K-|S|}{\sum_{k:(u^{\star}(X_{0}))_{k}=0}(\bar{A}_{t})_{k}}+1\right)\\ &\leq\sum_{t=1}^{T}\sum_{k:(u^{\star}(X_{0}))_{k}=0}(\bar{A}_{t})_{k}\ln\frac{e(K-|S|)T}{\sum_{t=1}^{T}\sum_{k:(u^{\star}(X_{0}))_{k}=0}(\bar{A}_{t})_{k}}\\ &\leq\ln((K-|S|)T)\sum_{t=1}^{T}\sum_{k:(u^{\star}(X_{0}))_{k}=0}(\bar{A}_{t})_{k},\end{split}$ (9)
@@
Substituting the above expression into equation 9, noticing that $\mathbb{E}[\pi_{t}(u^{\star}(X_{0}))|\mathcal{F}_{t-1}]=\mathbb{E}[\pi_{t}(u^{\star}(X_{t}))|\mathcal{F}_{t-1}]$ and that $|S|\leq m$ thus establishes the claim. ∎
@@
For ease of notation, let us denote $\kappa=\sqrt{Kd\ln T+K(\ln T)^{2}/\lambda_{\min}(\Sigma)}.$ The regret bound stated in (i) follows immediately from Lemma 2.2 combined with Lemma 2.4 as well as the observation that $\sum_{t=1}^{T}H(\bar{A}_{t}(X_{0}))\leq m\ln(K/m)T.$ We now show claim (ii). Note that in the case of $\sum_{t=1}^{T}\sum_{k:(u^{\star}(X_{0}))_{k}=0}(\bar{A}_{t})_{k}<e$, we have $\sum_{t=1}^{T}H(\bar{A}_{t}(X_{0}))\leq e\ln(e(K-m)T)+1/e$, in which case we already have the desired bound. We thus continue to consider the case when $\sum_{t=1}^{T}\sum_{k:(u^{\star}(X_{0}))_{k}=0}(\bar{A}_{t})_{k}\geq e$. Observe that due to the definition of suboptimality gap $\Delta_{A}(x)$ and the corruption budget $C$, it follows that
@@
where the first inequality holds thanks to Lemma 2.2 combined with Lemma 2.4 as well as the lower bound equation 10, the second inequality follows by Lemma 2.5. The third inequality holds by the observation $a\sqrt{x}-bx\leq a^{2}/(4b)$ for any nonnegative scalars $a,b,x$, and the last equality follows by choosing $\lambda=1.$